# DiffPF: Differentiable Particle Filtering with Generative Sampling via Conditional Diffusion Models

### v1

@@ III-C End-to-End Training
Training the denoising diffusion model involves maintaining a forward diffusion process and learning to reverse this process. Instead of explicitly iterating through each diffusion step, we utilize a closed-form solution to efficiently sample a noisy version of the ground truth state $\bm{x}_{t,0}^{*}$ at an arbitrary diffusion timestep $k$:
$\bm{x}_{t,k}^{*}=\sqrt{\bar{\alpha}}\bm{x}_{t,0}^{*}+\sqrt{1-\bar{\alpha}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,I),$ (11)
where $\bm{\epsilon}$ denotes a Gaussian noise. The model is then trained to denoise $\bm{x}_{t,k}^{*}$ by predicting the added noise $\bm{\epsilon}$, conditioned on the timestep $k$ and the fusion vector $\bm{c}_{t}$. A U-Net $\epsilon_{\theta}$ (see LABEL:fig:teaser) is used for this task, and training minimizes the mean squared error between predicted and actual noise:
$\mathcal{L}=\mathbb{E}_{\bm{x}_{t,k}^{*},k,{\bm{c}}_{t},\epsilon}\left[\left\|\bm{\epsilon}-\epsilon_{\theta}(\bm{x}_{t,k}^{*},k,{\bm{c}}_{t})\right\|^{2}\right].$ (12)
Each training iteration consists of the following procedure: (1) Draw a ground truth state $\bm{x}_{t,0}^{*}$ from the data; (2) Sample a timestep $k$ and generate its noisy version $\bm{x}^{*}_{t,k}$ using  (11); (3) Use $\epsilon_{\theta}$ to predict the noise from $\bm{x}_{t,k}^{*}$, $k$, and $\bm{c}_{t}$; (4) Update model parameters via minimizing the loss (12).
Importantly, all components of the system—including the process model, observation model, and fusion module—are trained end-to-end along with the diffusion model to ensure consistency in optimization.
@@ IV EXPERIMENTS
We evaluate DiffPF across three challenging state estimation tasks: (1) a synthetic image-based object tracking task, as introduced in [41, 22]; (2) a simulated global localization task with pronounced multimodality, based on [9, 19]; and (3) the real-world KITTI visual odometry benchmark [11].
We compare DiffPF against a comprehensive set of baselines in the following experiments, including the Deep State-Space Model (Deep SSM) [7], Particle Filter Recurrent Neural Networks (PFRNNs) [23], Particle Filter Networks (PFNets) [10], Normalizing Flow-based Differentiable Particle Filters (NF-DPFs) [18], Autoencoder Sequential Monte Carlo (AESMC) and its bootstrap variant (AESMC-Bootstrap) [24, 25], as well as other differentiable filtering and smoothing methods [22, 42]. Our methods were trained and evaluated on a workstation equipped with an NVIDIA GeForce RTX 4080 GPU and an Intel i7-14700KF CPU.
Experimental results demonstrate that DiffPF generalizes well across all test scenarios, achieves significantly higher state estimation accuracy compared to existing differentiable filters, and maintains practical inference efficiency.
@@ IV-A Vision-based Disk Tracking
Data: We divide the dataset into 500 sequences for training and 50 for testing, with each sequence consisting of 50 RGB images at a resolution of 128×128 pixels following [22]. Each image contains a target disk and 25 distractor disks. The target disk consistently appears in red with a fixed radius of 7 pixels while the remaining 25 distractor disks vary in both size and color, with radii randomly selected from 3 to 10 pixels, and colors sampled from a predefined palette same as [18]. All disks are initially placed uniformly across the frame. The dynamics of all disks follow
$\left\{\begin{array}[]{ll}\bm{x}^{*}_{t+1}=\bm{x}^{*}_{t}+\bm{v}_{t}^{*}+\bm{\omega}_{x,t}\\ \bm{v}_{t+1}^{*}=\bm{v}_{t}^{*}-f_{p}\bm{x}_{t}^{*}-f_{d}\|\bm{v}_{t}^{*}\|^{2}\,\mathrm{sgn}(\bm{v}_{t}^{*})+\bm{\omega}_{v,t},\end{array}\right.$ (13)
where $\bm{x}^{*}_{t}$ and $\bm{v}^{*}_{t}$ denote the position and velocity at time $t$; $f_{p}=0.05$ is a coefficient that pulls the object toward the origin; $f_{d}=0.0075$ controls the strength of velocity damping; $\bm{\omega}_{x,t}$ and $\bm{\omega}_{v,t}$ are Gaussian noise; and $\mathrm{sgn}(\cdot)$ returns the sign of a vector component-wise.
Implementation: The number of particles is set to 10. At time step $t$, we use the given velocity $\bm{v}_{t}^{*}$ as input to generate a prior estimate of $\bm{x}_{t-1}^{(i)}$. The following formulation serves as our process model:
$\hat{\bm{x}}_{t}^{(i)}=\bm{x}_{t-1}^{(i)}+\bm{v}^{*}$ (14)
Due to the imperfection of the process model, observation $\bm{o}_{t}$ is used to correct the prior estimate. We adopt the network architecture from [18] as our sensor model. To integrate $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ with $\bm{o}_{t}$, we first convert the $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ into a heatmap. The observation $\bm{o}_{t}$ is encoded by the sensor model, while the heatmap is processed by a separate network with the same architecture. The outputs of both networks are then concatenated to form the feature representation $\bm{c}_{t}$, which are used as conditions for the diffusion model. The noise estimator in the diffusion model is implemented as a 3-layer U-Net.
@@
1)))) Benchmarking. We compare the DiffPF with several baselines, including DeepSSM, AESMC-Bootstrap, AESMC, PFRNN, PFNet, and NF-DPF, all of which adopt the same process and observation models where applicable.
2)))) Ablation analysis. We further investigate the design of our method through controlled experiments, including:
@@
  • Introduction of the model prior. We also evaluate DiffPF(no pred), a variant with 10 diffusion steps that does not use the process model and relies only on observations.
3)))) Inference efficiency. We assess the runtime of the DiffPF under different diffusion step settings.
@@
Predicted State DiffPF (no pred) 5.23 $\pm$ 0.18
TABLE II: Runtime performance for different numbers of diffusion steps. Mean and standard deviation are computed from 5 independent runs with varying random seeds.
Diffusion Step K = 5 K = 10
Inference Frequency (Hz) 59.0 $\pm$ 2.75 39.2 $\pm$ 1.11
Comparison: As shown in Table II and Table I, we observe that:
1. DiffPF outperforms existing differentiable particle filters in both accuracy and consistency while requiring far fewer particles. Compared to NF-DPF, which also leverages generative modeling, it shows a 62.2 % improvement. Moreover, the significantly lower variance across runs highlights its robustness and reliability.
2. Using more diffusion steps improves estimation accuracy, suggesting that a longer denoising process enhances the model’s predictive capability. Moreover, incorporating the predicted state from the process model results in a 73.8% reduction in MSE, underscoring its substantial contribution to the overall performance.
3. Inference speed decreases with more diffusion steps. However, both settings sustain relatively high inference rates, demonstrating the practical applicability of DiffPF.
@@ IV-B Global Localization
Overview: The task is to estimate the robot’s pose based on observed images and odometry data, within three maze environments simulated in DeepMind Lab [43]. The main challenge of this task lies in the highly multimodal nature of the observations: since we adopt the same experimental setup as in [9], unique textures have been deliberately removed from the environment, causing many locations in the maze to produce nearly identical images. An example of the observation image is shown in Fig 3, along with the corresponding candidate poses on the map.
@@
Data: Consistent with [18], the dataset is divided into 900 training sequences and 100 test sequences, with each sequence containing 100 time steps. The ground truth state is defined as the robot’s position and orientation, denoted as $\bm{x}_{t}^{*}:=(p_{t,x}^{*},p_{t,y}^{*},\theta_{t}^{*})$, and the action $\bm{a}_{t}^{*}$ is computed from the state difference between consecutive time steps, $\Delta\bm{x}_{t}^{*}=\bm{x}_{t}^{*}-\bm{x}_{t-1}^{*}=(\Delta p_{t,x}^{*},\Delta p_{t,y}^{*},\Delta\theta_{t}^{*})$ as
$u_{t}^{*}$ $=\cos(\theta_{t-1}^{*})\cdot\Delta p_{t,x}^{*}+\sin(\theta_{t-1}^{*})\cdot\Delta p_{t,y}^{*},$ (15)
$v_{t}^{*}$ $=-\sin(\theta_{t-1}^{*})\cdot\Delta p_{t,x}^{*}+\cos(\theta_{t-1}^{*})\cdot\Delta p_{t,y}^{*},$
$\bm{a}_{t}^{*}$ $:=(u_{t}^{*},v_{t}^{*},\Delta\theta_{t}^{*}),$
which corresponds to the robot’s velocity in its local frame. The original $32\times 32$ RGB observations are randomly cropped to $24\times 24$ and augmented with Gaussian noise ($\sigma=20$) to introduce visual uncertainty.
Implementation: To effectively capture the uncertainty arising from the task’s multimodal structure, we use 20 particles in our filter. At time step $t$, we use the given input $\bm{a}_{t}^{*}$ to generate a prior estimate of $\bm{x}_{t}^{(i)}=(p_{t,x}^{(i)},p_{t,y}^{(i)},\theta_{t}^{(i)})$. We form the process model as:
@@
where $\bm{\omega}_{t}$ denotes zero-mean Gaussian noise with covariance $\mathbf{\Sigma}=\mathrm{diag}(10^{2},10^{2},0.1^{2})$. We adopt the same encoder from [18] as our sensor model. To integrate $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ with $\bm{o}_{t}$, We encode the observation $\bm{o}_{t}$ using the sensor model and flatten the particle set $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ into a vector. The two are concatenated to form the conditional input $\bm{c}_{t}$ for the diffusion model. The noise estimator in the diffusion model is implemented as a 3-layer U-Net. The model is first pretrained using only observations, and then finetuned in a second training phase with particle-based priors integrated.
Results: DiffPF is compared with the same set of baselines as in the previous benchmark under a shared setup to assess overall performance. To analyze the contribution of the process model prior, we include a variant trained without it denoted as DiffPF (no pred). With the diffusion step set to 10, the model achieves an inference frequency of 45.5 Hz. All methods are evaluated using the root mean squared error (RMSE) between the predicted and ground-truth positions of the robot at the last step. The results are summarized in Table III.
@@
DiffPF 12.3 $\pm$ 0.3 11.2 $\pm$ 0.3 25.7 $\pm$ 2.7
@@
Comparison: As shown in Table III, DiffPF achieves superior performance over existing differentiable particle filters in terms of both accuracy and consistency, while operating with substantially fewer particles. Compared to NF-DPF—which also employs generative modeling—DiffPF achieves improvements of 73.3%, 89.1%, and 85.9% in Maze 1, Maze 2, and Maze 3, respectively. Furthermore, the notably lower variance observed across multiple runs underscores its robustness and reliability. As shown in Fig. 3, even in the most complex and highly multimodal Maze 3 environment, DiffPF is able to rapidly converge to a pose close to the ground truth during the early stages, while other methods may require 50 or more steps to achieve similar accuracy, highlighting the strong representational and sampling capabilities of the diffusion model. Incorporating the process model prior yields RMSE reductions of 91.0%, 94.7%, and 93.8% in the three maze environments, indicating that introducing the prior effectively reduces the multimodal ambiguity in the observation-to-state mapping.
@@ IV-C KITTI Visual Odometry
Overview: We evaluate our method on the KITTI Visual Odometry dataset [11], where the objective is to track a vehicle’s pose over time using RGB images captured from a stereo camera and the given initial pose. In this task, the vehicle’s pose is not directly available from images and must be inferred by estimating relative motion across frames and accumulating it over time. This process is prone to drift, especially without loop closure. Moreover, the limited size and diversity of the dataset pose challenges for generalization to unseen road conditions.
Data: Following the protocol from [22], we adopt the KITTI-10 dataset, which comprises ten driving sequences recorded in urban environments. Each sequence includes paired stereo images and ground truth poses. The data is recorded at around 10 Hz. To expand the training data, we perform horizontal flipping on the stereo image pairs. After augmentation, sequences are segmented into fixed-length clips and downsampled to 50$\times$150 pixels per frame.
Filter Implementation: The vehicle’s state at each time step is represented as a 5D vector $\bm{x}_{t}=(l^{x}_{t},l^{y}_{t},\theta_{t},v_{t},\omega_{t})$, consistent with [22, 42]. Here, $(l^{x}_{t},l^{y}_{t})$ denote the 2D position, $\theta_{t}$ is the heading angle, $v_{t}$ is the linear velocity, and $\omega_{t}$ the angular rate. Observations include the current RGB frame and a difference frame computed by subtracting the previous frame, following [41] (see Fig. 4). We use 10 particles in our filter.
Figure 4: Each row shows an example from a different sequence: (Left) frame at $t{-}1$, (Middle) current frame, and (Right) the corresponding difference frame.
Since explicit control inputs are not available, we adopt a constant velocity and angular velocity model as the process model, where the change in velocities between consecutive time steps is treated as process noise. The sensor model for observed images reuses the backbone architecture from [22, 42]. We encode the observation $\bm{o}_{t}$ using the sensor model and flatten the predicted particle set $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ into a vector. The two are concatenated to form the conditional input $\bm{c}_{t}$ for the diffusion model. The denoising network in diffusion model is instantiated as a same 3-layer U-Net. The training employs a 10-fold cross-validation setup, where nine trajectories are used for training and the remaining one for testing in each fold. The model is pretrained solely based on visual observations, the particle set is introduced in a subsequent training stage.
Results: We use 10 diffusion steps and evaluate on 100-frame segments following KITTI’s standard metrics. Positional and angular errors are computed by comparing the final predicted pose to the ground truth and normalizing by the total traveled distance (m/m and deg/m, respectively).
@@
1) Differentiable Filters: Include several state-of-the-art differentiable recursive filters proposed in previous studies [22, 41, 35]. To ensure a fair evaluation, we adopt the same input and use identical sensor model architectures.
2) LSTMs: Building on previous research in differentiable filtering [22, 41], we include long short-term memory (LSTM) models [44]. We choose an unidirectional LSTM and a LSTM integrated with dynamics as the baselines [22].
3) Smoothers: We adopt the differentiable smoothers introduced in [42], which are built upon factor graph optimization. As these methods utilize the full observation sequence, they naturally hold an advantage over filtering-based approaches.
All of the experimental results are presented in Table IV.
TABLE IV: Performance comparison of DiffPF and baseline models on KITTI dataset.
@@
Comparison: Our DiffPF achieves a 25% reduction in position error and a 26% reduction in orientation error compared to leading differentiable filtering approaches. It also surpasses LSTMs and differentiable smoothers in performance. These results highlight the strength of the diffusion model in capturing complex posterior distributions, enabling more effective particle sampling and higher-quality approximations of the posterior. Furthermore, the method operates at 46.5 Hz, well above the 10 Hz input rate of the visual stream, ensuring its suitability for real-time applications.

### v2

@@ III-C End-to-End Training
The diffusion model is trained following the standard DDPM formulation  [ho2020denoising]. At each training iteration, a noisy version of the ground-truth state $\bm{x}_{t,k}^{*}$ is generated by adding Gaussian noise to the clean state $\bm{x}_{t,0}^{*}$, based on a randomly selected diffusion timestep $k$. The denoising network $\epsilon_{\theta}$, implemented as a U-Net (see Figure˜1), is trained to predict the added noise $\bm{\epsilon}$ conditioned on the diffusion timestep $k$ and the fusion vector $\bm{c}_{t}$. The training objective minimizes the mean squared error between the predicted and true noise:
$\mathcal{L}=\mathbb{E}_{\bm{x}_{t,0}^{*},k,\bm{c}_{t},\bm{\epsilon}}\left[\|\bm{\epsilon}-\epsilon_{\theta}(\bm{x}_{t,k}^{*},k,\bm{c}_{t})\|^{2}\right].$ (9)
All components of the system—including the process model, observation encoder, and fusion module—are trained jointly with the diffusion model in an end-to-end manner to ensure consistent optimization across the entire framework.
@@ IV EXPERIMENTS
We evaluate DiffPF across four challenging state estimation tasks: (1) a synthetic image-based object tracking task; (2) a simulated global localization task with pronounced multimodality; (3) the real-world KITTI visual odometry benchmark and (4) a robotic manipulation task.
We compare DiffPF against a comprehensive set of baselines in the following experiments, including the Deep State-Space Model (Deep SSM) [rangapuram2018deep], Particle Filter Recurrent Neural Networks (PFRNNs) [ma2020particle], Particle Filter Networks (PFNets) [karkus2018particle], Normalizing Flow-based Differentiable Particle Filters (NF-DPFs) [chen2024normalizing], Autoencoder Sequential Monte Carlo (AESMC) and its bootstrap variant (AESMC-Bootstrap) [le2018auto, naesseth2018variational], as well as other differentiable filtering and smoothing methods [kloss2021train, yi2021differentiable, liu2023alpha, liu2023enhancing]. All DPF-based baselines are evaluated with 100 particles, while DiffPF uses 10 particles unless otherwise specified. The noise estimator in the diffusion model is implemented as a 3-layer U-Net in all experiments and the number of diffusion steps is set to 10 unless otherwise specified. DiffPF was trained and evaluated on a workstation equipped with an NVIDIA GeForce RTX 4080 GPU and an Intel i7-14700KF CPU. Training uses the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$ and a cosine annealing schedule. We train for 1000 epochs with warmup enabled for the first 4 epochs. For the disk tracking task we use a batch size of 50, while for global localization the batch size is 100. For KITTI visual odometry and robotic manipulation task, due to the fold-based training protocol, the batch size varies across folds. Full configurations are provided in our released code.
@@ IV-A Vision-based Disk Tracking
Data: We follow exactly the same dataset and simulation setup as in [chen2024normalizing]. The dataset consists of 500 training and 50 testing sequences, each containing 50 RGB frames of size 128×128. Each frame includes one target disk and 25 distractor disks, initialized uniformly at random. The target red disk has a fixed radius of 7 pixels, while distractors vary in size and color. The disk dynamics follow the same formulation as in [kloss2021train].
Implementation: At time step $t$, we use the given velocity $\bm{v}_{t}^{*}$ as input to generate a prior estimate of $\bm{x}_{t-1}^{(i)}$. The following formulation serves as our process model:
$\hat{\bm{x}}_{t}^{(i)}=\bm{x}_{t-1}^{(i)}+\bm{v}_{t}^{*}.$ (10)
Due to the imperfection of the process model, observation $\bm{o}_{t}$ is used to correct the prior estimate. We adopt the network architecture from [chen2024normalizing] as our sensor model. To integrate $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ with $\bm{o}_{t}$, we first convert the $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ into a heatmap. The observation $\bm{o}_{t}$ is encoded by the sensor model, while the heatmap is processed by a separate network with the same architecture. The outputs of both networks are then concatenated to form the feature representation $\bm{c}_{t}$, which are used as conditions for the diffusion model.
@@
1)) Benchmarking. We compare the DiffPF with several baselines, including DeepSSM, AESMC-Bootstrap, AESMC, PFRNN, PFNet, and NF-DPF, all of which adopt the same process and observation models where applicable.
2)) Ablation analysis. We further investigate the design of our method through controlled experiments, including:
@@
  • Effect of the process model prior. We perform an ablation study by injecting different levels of noise into the process model Eq. 10. The variants are denoted as DiffPF($\sigma$), where $\sigma$ controls the standard deviation of the noise added to the dynamics. We also test the variant that removes the prior, denoted as DiffPF(no pred).
3)) Inference efficiency. We assess the runtime of the DiffPF under different diffusion step settings.
@@
Priors Accuracy DiffPF (no pred) 5.23 $\pm$ 0.18
DiffPF ($\sigma=2$) 1.44 $\pm$ 0.01
DiffPF ($\sigma=5$) 1.96 $\pm$ 0.02
DiffPF ($\sigma=10$) 3.37 $\pm$ 0.12
TABLE II: Inference frequency for different numbers of diffusion steps. Means are computed from 5 independent runs with varying random seeds.
Diffusion Step K = 5 K = 10 K = 20 K = 50 K = 100
Infer. Freq. (Hz) 59.0 39.2 21.1 9.9 5.26
Comparison: As shown in Table I and Table II, DiffPF consistently outperforms existing differentiable particle filters in terms of both accuracy and stability while using significantly fewer particles, achieving a 62.2% improvement over NF-DPF. Increasing the number of diffusion steps further improves estimation accuracy, while inference remains real-time in most settings with a modest GPU memory footprint of 224.1 MB. Incorporating the predicted state from the process model yields a 73.8% reduction in MSE and DiffPF maintains strong robustness under prior perturbations.
@@ IV-B Global Localization
Overview: The task is to estimate the robot’s pose based on observed images and odometry data, within three maze environments simulated in DeepMind Lab [beattie2016deepmind]. The main challenge of this task lies in the highly multimodal nature of the observations: since we adopt the same experimental setup as in [jonschkowski2018differentiable], unique textures have been deliberately removed from the environment, causing many locations in the maze to produce nearly identical images. An example of the observation image is shown in Fig 4, along with the corresponding candidate poses on the map.
@@
Data: Consistent with [chen2024normalizing], the dataset is divided into 900 training sequences and 100 test sequences, with each sequence containing 100 time steps. The ground truth state is defined as the robot’s position and orientation, denoted as $\bm{x}_{t}^{*}:=(p_{t,x}^{*},p_{t,y}^{*},\theta_{t}^{*})$, and the action $\bm{a}_{t}^{*}=(u^{*}_{t},v_{t}^{*},\Delta\theta^{*}_{t})$ corresponds to the robot’s velocity in its local frame. The original $32\times 32$ RGB observations are randomly cropped to $24\times 24$ and augmented with Gaussian noise ($\sigma=20$) to introduce visual uncertainty.
Implementation: At time step $t$, we use the given input $\bm{a}_{t}^{*}$ to generate a prior estimate of $\bm{x}_{t}^{(i)}=(p_{t,x}^{(i)},p_{t,y}^{(i)},\theta_{t}^{(i)})$. We form the process model as:
@@
where $\bm{\omega}_{t}$ denotes zero-mean Gaussian noise with covariance $\mathbf{\Sigma}=\mathrm{diag}(10^{2},10^{2},0.1^{2})$. We adopt the same encoder from [chen2024normalizing] as our sensor model. To integrate $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ with $\bm{o}_{t}$, We encode the observation $\bm{o}_{t}$ using the sensor model and flatten the particle set $\{\hat{\bm{x}}_{t}^{(i)}\}_{i=1}^{N}$ into a vector. The two are concatenated to form the conditional input $\bm{c}_{t}$ for the diffusion model. The model is first pretrained using only observations, and then finetuned in a second training phase with particle-based priors integrated.
Results: DiffPF is compared with the same set of baselines as in the previous benchmark under a shared setup to assess overall performance. To further analyze the framework, we conduct several ablation studies. First, to evaluate the contribution of the process model prior, we include a variant trained without it, denoted as DiffPF (no pred), and additional variants where different levels of noise are injected into the prior (DiffPF($\Sigma$)). Second, we investigate particle efficiency by varying the number of particles and measuring both inference frequency and memory footprint across tasks. All methods are evaluated using the root mean squared error (RMSE) between the predicted and ground-truth positions of the robot at the last step. The results are summarized in Table III,  IV and  V.
@@
DiffPF 6.1 $\pm$ 0.5 7.6 $\pm$ 0.2 15.3 $\pm$ 1.2
DiffPF($2\Sigma$) 8.7 $\pm$ 0.3 11.7 $\pm$ 0.5 17.8 $\pm$ 1.3
DiffPF($5\Sigma$) 18.7 $\pm$ 1.2 25.2 $\pm$ 2.5 28.4 $\pm$ 2.4
DiffPF($10\Sigma$) 106.4 $\pm$ 16.1 61.7 $\pm$ 9.2 67.2 $\pm$ 17.4
@@
TABLE IV: Effect of particle number NN on RMSE at the last step $t=100$, averaged over 5 runs.
Particle Number Maze 1 Maze 2 Maze 3
N = 1 19.3 $\pm$ 14.2 37.3 $\pm$ 27.6 131.3 $\pm$ 53.0
N = 5 13.2 $\pm$ 0.9 11.1 $\pm$ 2.5 30.9 $\pm$ 14.2
N = 10 6.1 $\pm$ 0.5 7.6 $\pm$ 0.2 15.3 $\pm$ 1.2
N = 20 5.4 $\pm$ 0.2 7.6 $\pm$ 0.3 25.7 $\pm$ 2.7
N = 40 4.9 $\pm$ 0.1 8.2 $\pm$ 0.1 20.1 $\pm$ 6.8
N = 80 4.3 $\pm$ 0.2 7.7 $\pm$ 0.1 17.2 $\pm$ 0.4
TABLE V: Inference frequency and memory footprint for different numbers of particles. Means are computed from 5 independent runs with varying random seeds.
Particle number N = 5 N = 10 N = 20 N = 40 N = 80
Infer. Freq. (Hz) 54.1 52.6 45.5 27.4 14.3
Memory (MB) 188.3 194.2 205.5 230.0 276.6
Comparison: As shown in Table III, DiffPF achieves superior performance over existing differentiable particle filters in terms of both accuracy and consistency, while operating with substantially fewer particles. Compared to NF-DPF—which also employs generative modeling—DiffPF achieves improvements of 86.8%, 92.6%, and 91.6% in Maze 1, Maze 2, and Maze 3, respectively. Furthermore, the notably lower variance observed across multiple runs underscores its robustness and reliability.
@@
As shown in Fig. 5, even in the most complex and highly multimodal Maze 3 environment, DiffPF is able to rapidly converge to a pose close to the ground truth during the early stages, while other methods may require 50 or more steps to achieve similar accuracy. Incorporating the process model prior yields RMSE reductions of 95.5%, 96.4%, and 96.3% in the three maze environments, indicating that introducing the prior effectively reduces the multimodal ambiguity in the observation-to-state mapping. Furthermore, DiffPF demonstrates strong robustness under moderate perturbations of the prior, exhibiting only gradual performance degradation. However, when the noise level becomes extremely large (e.g., $10\Sigma$), as shown in Fig. 6, prior inaccuracies can occasionally lead to sample collapse, manifested as convergence to incorrect trajectories. This phenomenon likely stems from multiple factors, including limited coverage of such extreme perturbations in the training data, insufficient convergence due to suboptimal training or optimization, and constrained model capacity under highly perturbed dynamics. Nevertheless, even under this challenging setting, DiffPF still outperforms almost all baseline methods.
Figure 6: An example where the proposed DiffPF collapses under highly noisy priors. The arrow notation is the same as in Fig. 5.
As for the effect of particle number, Table IV shows that increasing $N$ generally improves estimation accuracy and stability. Notably, DiffPF already achieves low RMSE with as few as $N=10$ particles, and further gains beyond $N=40$ are relatively modest. This demonstrates the particle efficiency of our approach compared to traditional filters, which require more particles (100 particles) yet still fail to reach similar accuracy. In terms of scalability, inference frequency decreases and memory usage increases approximately linearly with $N$. Nevertheless, even with $N=80$, DiffPF sustains real-time inference rates with moderate memory consumption, confirming its practicality while delivering strong accuracy with relatively few particles.
@@ IV-C KITTI Visual Odometry
Overview: We evaluate our method on the KITTI Visual Odometry dataset [geiger2013vision], where the objective is to track a vehicle’s pose over time using RGB images captured from a stereo camera and the given initial pose. In this task, the vehicle’s pose is not directly available from images and must be inferred by estimating relative motion across frames and accumulating it over time. This process is prone to drift, especially without loop closure. Moreover, the limited size and diversity of the dataset pose challenges for generalization to unseen road conditions.
Data: Following the protocol in [kloss2021train, wan2025dnd], we evaluate DiffPF on the KITTI-10 dataset, which consists of ten urban driving sequences with paired stereo images and ground-truth poses recorded at approximately 10 Hz. Stereo image pairs are horizontally flipped for data augmentation, and the sequences are segmented into fixed-length clips and downsampled to $50\times 150$ pixels.
Implementation: The vehicle state is represented as a 5D vector and observations include the current RGB frame and a difference frame computed from the previous frame [kloss2021train, haarnoja2016backprop]. The process model design, sensor model network, and denoising network follow those in [wan2025dnd]. The predicted particle set is first flattened and concatenated with the observation features to form the conditioning input for the diffusion model. Training follows a 10-fold cross-validation protocol, with the model pretrained using only visual observations before introducing the particle set.
Results: We evaluate on 100-frame segments following KITTI’s standard metrics. Positional and angular errors are computed by comparing the final predicted pose to the ground truth and normalizing by the total traveled distance (m/m and deg/m, respectively).
@@
1) Differentiable Filters: Include several state-of-the-art differentiable recursive filters proposed in previous studies [kloss2021train, haarnoja2016backprop, wan2025dnd]. To ensure a fair evaluation, we adopt the same input and use identical sensor model architectures.
2) LSTMs: Building on previous research in differentiable filtering [kloss2021train, haarnoja2016backprop], we include long short-term memory (LSTM) models [graves2012long]. We choose an unidirectional LSTM and a LSTM integrated with dynamics as the baselines [kloss2021train].
3) Smoothers: We adopt the differentiable smoothers introduced in [yi2021differentiable], which are built upon factor graph optimization. As these methods utilize the full observation sequence, they naturally hold an advantage over filtering-based approaches.
All of the experimental results are presented in Table VI.
TABLE VI: Performance comparison of DiffPF and baseline models on KITTI dataset.
@@
Comparison: Our DiffPF achieves a 25% reduction in position error and a 26% reduction in orientation error compared to leading differentiable filtering approaches. It also surpasses LSTMs and differentiable smoothers in performance. These results highlight the strength of the diffusion model in capturing complex posterior distributions, enabling more effective particle sampling and higher-quality approximations of the posterior. Furthermore, the method operates at 46.5 Hz with a memory footprint of 222.7 MB, well above the 10 Hz input rate of the visual stream, ensuring its suitability for real-time applications.
IV-D Robotic Manipulation
Overview: We evaluate our method on a robot manipulation state estimation task described in [liu2023alpha]. The objective is to track the internal state of a UR5 robot during tabletop manipulation from RGB image sequences, as illustrated in Fig. 7. Under partial observability, the system must estimate and propagate a higher-dimensional robot state purely from visual inputs. Experiments are conducted in both simulated and real-world environments to assess robustness across domains.
Data: Following the protocol and dataset of [liu2023alpha], we consider UR5 manipulation sequences collected in simulation and on a real robot. The internal state $x\in\mathbb{R}^{10}$ composed of 7 joint angles and the Cartesian position of the end-effector. RGB images are used as observation. All sequences are segmented into fixed-length clips for training and evaluation.
Implementation: Since actions are not available and the data are sampled at a relatively high frequency, the process model adopts an identity transition, setting $\hat{x}_{t}^{(i)}=x_{t-1}^{(i)}$. The sensor model network follows the design in [liu2023alpha]. To stabilize training, we first pretrain the model using visual observations only.
Results: We evaluate performance using mean absolute error (MAE) on joint angles and end-effector positions. All results are reported in comparison with the methods and baselines presented in [liu2023alpha, liu2023enhancing] and presented in Table VII.
Figure 7: Example image sequence from the UR5 manipulation task
TABLE VII: Result evaluations on UR5 manipulation task
Method Real-world (MAE) Simulation (MAE)
Joint (deg) EE (cm) Joint (deg) EE (cm)
[…]
Means $\pm$ standard errors.
Comparison: DiffPF achieves substantially lower estimation errors than existing differentiable filtering methods on the UR5 manipulation task. In real-world experiments, it reduces end-effector position error by 48% and joint angle error by 43% compared to the strongest baseline, while exhibiting consistent improvements in simulation. In addition, DiffPF operates at over 50 Hz, demonstrating its real-time applicability. These results further indicate that DiffPF maintains strong performance even when scaling to higher-dimensional state spaces.